5. Show and explain the process of solving the following application problems using equations:
a. A cable 200 ft. long is cut into two pieces so that one piece is 12 ft. longer than the other. How long is each piece?
b. A boat travels 24 miles upstream in 3 hours. It then takes 3 hours to go 36 miles downstream. Find the speed of the boat in still water and the speed of the current.
c. A service station wishes to mix 40L of a 78% antifreeze solution. How many liters of a 75% solution and a 90% solution should be used in forming the mixture?

Answer by xcentaur(357)

a. A cable 200 ft. long is cut into two pieces so that one piece is 12 ft. longer than the other. How long is each piece?
let the first pc be x ft long
then second piece is (x+12)
together they make up 200 feet
so x+x+12=200
2x+12=200
2x=200-12=188
x=188/2=94


b. A boat travels 24 miles upstream in 3 hours. It then takes 3 hours to go 36 miles downstream. Find the speed of the boat in still water and the speed of the current.
let speed of boat in still water be x,and speed of current downstream be y
then upstream the speed of the boat will be (y-x) miles per hour
distance=24 miles in 3 hours at speed of (y-x)
speed=(y-x)=distance/time=24/3=8
speed downstream is (x+y)
speed=(x+y)=dstnce/time=36/3=12
so we have 2 equations and 2 variables,
y-x=8 AND x+y=12
y=8+x AND y=12-x
y=8+x=12-x
8+x=12-x
x+x=12-8
2x=4
x=2
so speed of boat is 2 mph
y=8+x=10 mph which is speed of current


c. A service station wishes to mix 40L of a 78% antifreeze solution. How many liters of a 75% solution and a 90% solution should be used in forming the mixture?
let x litres be taken of the 75% solution
let y litres be taken of the 90% solution
then (x+y)=40 as we need 40litres of the solution
so y=40-x
now we can write our equation as,
x(75%)+y(90%)=40(78%)
x(75/100)+y(90/100)=40(78/100)
75x+90y=40(78)
75x+90y=3120
75x+90(40-x)=3120
75x+90(40)-90x=3120
-15x+3600=3120
-15x=3120-3600
-15x=-480
x=32
y=40-x=8
so you need 32 litres of 75% solution,and 8liters of 90% solution


Hope this helps,
-xC